Superior Markets, Inc., operates three stores in a large metropolitan area. A segmented absorption costing income statement fa. The breakdown of the selling and administrative expenses that are shown above is as follows: Total North Store South Storeb. The lease on the building housing the North Store can be broken with no penalty. c. The fixtures being used in the North SRequired 1 Required 2 Required 3 Required 4 Required 5 How much employee salaries will the company avoid if it closes the NorRequired 1 Required 2 Required 3 Required 4 Required 5 How much employment taxes will the company avoid if it closes the NortRequired 1 Required 2 Required 3 Required 4 Required 5 What is the financial advantage (disadvantage) of closing the North StRequired 1 Required 2 Required 3 Required 4 Required 5 Assume that the North Stores floor space cant be subleased. However,
Answer #1

1.Salary of new employee = $11,600

Delivery crew = $5,000

Sales Salaries = $55,200

Store manager Salaries = $26,000- 12,600 = $13,400

General Office Salaries = $6,300

Employee Salaries Avoided per Quarter = $91,500

2.Employment Taxes Avoided = 91,500*15% = $13,725

3.Financial Advantage = costs avoided –margin lost

= Direct Advertising cost + Store Rent +2/3rd Insurance +Utilities + Salaries Avoided + Taxes Avoided- 345,000

= 61,000 + 95,000 + 10,500*2/3 + 30,630 + 91,500 + 13,725 – 345,000

= -$46,145 i.e. disadvantage

4.North store should not be closed, as there is a loss in closing the store

5.Financial Advantage of closing the north store = Gross Margin from Increased sales of East Store – Disadvantage calculated above

= (702,000/1,560,000)*840,000*1/4 – 46,145

=$48,355

Since there is net advantage, the store should be closed.